### v5

# The Principle of Uncertain Maximum Entropy

Abstract
The Principle of Maximum Entropy is a rigorous technique for estimating an unknown distribution given partial information while simultaneously minimizing bias. However, an important requirement for applying the principle is that the available information be provided error-free (Jaynes, 1982). We relax this requirement using a memoryless communication channel as a framework to derive a new, more general principle. We show our new principle provides an upper bound on the entropy of the unknown distribution and the amount of information lost due to the use of a given communications channel is unknown unless the unknown distribution’s entropy is also known. Using our new principle we provide a new interpretation of the classic principle and experimentally show its performance relative to the classic principle and other generally applicable solutions. Finally, we present a simple algorithm for solving our new principle and an approximation useful when samples are limited.
Keywords: maximum entropy, uncertainty, information theory, information geometry
1 Introduction
We consider the problem of estimating an unknown probability distribution given samples taken from it that, due to noise or other mechanisms, do not provide a consistent estimation of the original distribution. We use a common memoryless communication channel model to ground our problem, in which a sender encodes a message sampled from the unknown distribution and transmits it over a possibly noisy channel to a receiver. However, we may have additional information about the unknown distribution in addition to the received samples that we also wish to incorporate into the estimate. Specifically, the structure of the unknown distribution may be known up to a set of parameters, reducing the problem to estimating the parameters of a so-structured distribution given the samples. To this end, we wish to employ the Principle of Maximum Entropy to find parameters producing a satisfying distribution with the least bias possible.
The Principle of Maximum Entropy states that among all alternative distributions that equally match available information the best choice is the one with the maximum entropy. As formalized by Jaynes (1957) and Shannon (1948) and commonly encountered in practice in various disciplines the information takes the form of constraints on expectations over feature functions of the distribution variable. The feature functions determine the structure of the estimated distribution and this configuration underpins a number of well known distributions that have the maximum entropy given their parameters, including the Gaussian and Gamma (Cover and Thomas, 2006).
However, an important assumption for the application of the principle to real-world scenarios is that the empirical feature expectations being constrained to are known (Jaynes, 1982). Or more restrictively requiring that there is only one possible empirical distribution that matches given informatino. But, this requirement may be impossible to satisfy in domains where significant noise or otherwise uncertainty exists surrounding the samples, though special cases do exist (Jaynes, 1982). Previous attempts to address this problem assume the uncertainty is due to hidden variables (Wang et al., 2012; Bogert and Doshi, 2022; Bogert et al., 2016; Hunsop Hong and Schonfeld, 2008) and use an expectation derived from the available error-free information to fill in the missing portion. We instead seek a more generally applicable principle that retains the desirable properties of the classic principle while uncertainty exists as to the empirical distribution as a result of the communication channel in use.
Our contributions are as follows,
  • We derive our new principle as an application of the classic principle to a noisy communication channel
  • We present our new principle as a hierarchical convex program and an algorithm to solve it
  • Through analysis we show our principle generalizes earlier principles and interpret the classic principle in light of our new one
  • We show our new principle produces bounds on the entropy of the unknown distribution and the amount of information lost when a given noisy communication channel is in use
  • We present a number of experimental results in randomly generated problems that demonstrate the performance of our new principle as well as an approximation useful when the provided samples are limited
2 Background
In this work, we will use the following conventions and notation. Random variables are indicated with an uppercase letter such as $W$ with the corresponding lowercase letter $w$ giving an individual state the variable may be in from a set of all possible values indicated in script $\mathcal{W}$. A probability distribution for $W$ is given by $P(W)$, with each probability $1\geq Pr(W=w)=Pr(w)\geq 0$. We use a subscript to indicate parameters of a distribution $P_{\lambda,\phi}(W)$ with ”ground truth” distributions indicated with a 0 ($P_{0}(W),Pr_{0}(w)$). Empirical distributions are indicated with a tilde diacritic $\tilde{Pr}(w)$, and may be defined in terms of sampling; for example, suppose that we produce $N$ i.i.d. samples from $P_{0}(W)$ and collect them into a multiset $\mathcal{S}$. We may compute the resulting empirical distribution as $\tilde{Pr}(w)\triangleq{\frac{1}{N}}\sum\nolimits_{w^{\prime}\in\mathcal{S}}\delta(w,w^{\prime})\penalty 10000\ \penalty 10000\ \forall\penalty 10000\ w\in\mathcal{W}$, where $\delta$ is the Kronecker delta. Finally, algorithm names will be denoted with TeleType.
2.1 Communications Channel Model
To accurately model the process of parameter estimation in the presence of noise, we use a simple discrete memoryless communications channel.
Figure 1: A discrete, memoryless communications channel showing the process of transmitting a message $w$ (top) and receiving (bottom)
Suppose a message $w$ is to be sent to a receiver. We model each $w$ as being sampled i.i.d. from a distribution $P_{0}(W)$, which is unknown to the receiver, with the distribution of all the sent messages $\tilde{P}(W)$. To transmit the message, it must first be physically encoded in the world using some method. In the communications domain this method is commonly a function of $w$ to aid in decoding by the receiver, here we model this method using the distribution $P(X|W)$ and assume this distribution is known to the receiver. This change is to allow for the possibility of messages from non-engineered or partially engineered sources in which $w$ only partially determines the encoding $x$.
After encoding, $x$ is sent over some channel to the receiver, where it is received as the signal $y$ after possibly being transformed by the channel. This process is modeled with the distribution $P(Y|X)$, which is known by the receiver. In the domain of communications, the receiver desires an estimate of the message, $\hat{w}$, which is determined by first processing $y$ to estimate $\hat{x}$ using $P(X|Y)$, and then decoding $\hat{x}$ into $\hat{w}$ using $P(W|X)$. By contrast in distribution estimation the receiver desires an estimate of the parameters of $P_{0}(W)$ using the distribution of all received signals $\tilde{P}(Y)$.
A foundational result in information theory (Shannon, 1948) shows that it is possible to choose an encoding function such that the decoding error is vanishingly small. As an example, suppose we have a discrete $\mathcal{W}$, an encoding that maps all $w$ to equally spaced points along the real number line, and a channel that adds zero-mean Gaussian noise to $x$. We can therefore choose the distance between each $x$ such that it is large compared to the standard deviation of the channel noise. Thus, a large enough distance will ensure the estimate $\hat{x}$ for each $y$ has arbitrarily small error, and thus we may also estimate $\tilde{P}(W)$ with an arbitrary small amount of error.
2.2 Principle of Maximum Entropy
Using $\tilde{P}(W)$, we may now proceed to estimate the parameters of $P_{0}(W)$. However, we may also wish to incorporate any extra information we have about the structure of $P_{0}(W)$. To this end, we define $K$ feature functions $\phi_{k}\colon\mathcal{W}\to\mathbb{R}$ that represent the relevant information we have about each $w$. We assume that $\phi$ is complete, correct and known to the receiver. We now require that our estimated distribution is constrained to match the expected features of the empirical distribution, $\Phi=\tilde{\Phi}$, or:
$\sum\limits_{w\in\mathcal{W}}Pr(w)\phi_{k}(w)=\sum\limits_{w\in\mathcal{W}}\tilde{Pr}(w)\phi_{k}(w)$ (1)
Unfortunately, however, there is a serious issue with this approach: the distribution $P(W)$ may be underconstrained, possibly leading to an infinite set of feasible $P(W)$ for any given $\tilde{\Phi}$. To resolve this the Principle of Maximum Entropy states that we should choose the one distribution among the feasible set with the maximum entropy, ensuring it contains the least information of all alternatives and thus is expected to be the least biased. The Principle of Maximum Entropy with feature expectation constraints is expressed as the following convex non-linear program:
[…]
In the event a prior distribution about $P_{0}(W)$ is given, we may use the Principle of Minimum Cross Entropy (Shore and Johnson, 1980) to replace the above objective with $min\penalty 10000\ D_{KL}(P(W)\mathrel{\|}\Pi(W))\penalty 10000\ =\penalty 10000\$$\sum\nolimits_{w\in\mathcal{W}}Pr(w)\log{\frac{Pr(w)}{\Pi(w)}}$, where $D_{KL}$ is the Kullback-Leibler divergence and $\Pi(W)$ is the prior. By setting $\Pi(W)$ to the uniform distribution we arrive at the Principle of Maximum Entropy again. The solution to this program is the well known Gibbs distribution,
$Pr_{\lambda,\phi}(w)=\frac{e^{\sum\nolimits_{k}\lambda_{k}\phi_{k}(w)}}{Z_{\phi}(\lambda)}$ (3)
where $\lambda$ are the parameters of the Lagrangian dual of equation 2 and $Z_{\phi}$ is the partition function. Returning to parameter estimation, the receiver’s task is reduced to choosing $\lambda$ such that the one maximum entropy distribution which matches the feature expectations of $\tilde{P}(W)$ is found.
3 Uncertain $\tilde{P}(W)$
Unfortunately, in many scenarios in which we wish to apply the Principle of Maximum Entropy it is not possible to choose an encoding such that there is arbitrarily low estimation error per $w$ therefore $\tilde{P}(W)$ will remain uncertain. However, we may still apply the principle provided $\tilde{\Phi}$ is a consistent estimator. For instance, if $\Phi$ is the mean of $P_{0}(W)$, the sample mean $\tilde{\Phi}$ is a consistent estimator as it approaches $\Phi$ as $N\rightarrow\infty$.
In general, however, uncertainty as to $\tilde{P}(W)$ results in unavoidable uncertainty about $\tilde{\Phi}$, preventing the use of the Principle of Maximum Entropy unless we relax our constraints appropriately (Jaynes, 1982; Skilling and Bryan, 1984) or choose a single $\tilde{\Phi}$ among those possible. Toward the later approach we precisely define the possible $\tilde{\Phi}$; in this section we ignore the relationship between $\tilde{P}(W)$ and all $P_{\lambda,\phi}(W)$ and focus entirely on what is known about the empirical distribution. Given an arbitrarily accurate estimate of $\tilde{P}(Y)$, we have:
$\sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\sum\nolimits_{x\in\mathcal{X}}Pr(x|w)Pr(y|x)=\tilde{Pr}(y)\penalty 10000\ \penalty 10000\ \forall y$ (4)
We may now employ linear algebra to determine all satisfying $\tilde{P}(W)$ as follows. Gather all $\tilde{P}(W)$ into the vector $\mathbb{W}$, $\tilde{P}(Y)$ into the vector $\mathbb{Y}$, and define the $|\mathcal{Y}|x|\mathcal{W}|$ matrix $A$ as $A_{y,w}=\sum\nolimits_{x\in\mathcal{X}}Pr(x|w)Pr(y|x)$. Now: $A\mathbb{W}^{T}=\mathbb{Y}^{T}$. We can readily see that the null space of matrix A (and any one satisfying $\tilde{P}(W)$) defines the set of all $\tilde{P}(W)$ that may have ultimately produced $\tilde{P}(Y)$. When the null space is non-empty, such as when $|\mathcal{Y}|<|\mathcal{W}|$, we have a linear space of $\tilde{P}(W)$ any of which may equally have produced $\tilde{P}(Y)$.
In this situation existing methods could be used to choose one $\tilde{P}(W)$ so that we may then employ the Principle of Maximum Entropy. As some examples, we could employ Baye’s law and choose the most likely $\tilde{P}(W)$ given a prior on $P_{0}(W)$ by solving:
$\tilde{Pr}(w)=\sum\limits_{y}\tilde{Pr}(y)\penalty 10000\ \delta(w,\mathop{\mathrm{argmax}}\limits_{w^{\prime}}P(W=w^{\prime}|Y=y))$ (5)
Where $\delta$ is the Kronecker delta. However, while a simple solution, we are not guaranteed that the found $\tilde{P}(W)$ will satisfy equation 4 and this would ensure $\tilde{P}(W)$ could not be $P_{0}(W)$. Instead, we could employ on the Principle of Maximum Entropy to arrive at the $\tilde{P}(W)$ with the maximum entropy that satisfies equation 4.
$\begin{array}[]{l}\max-\sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\log\tilde{Pr}(w)\\ \mbox{{\bf subject to}}\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)=1\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\sum\nolimits_{x\in\mathcal{X}}Pr(x|w)Pr(y|x)=\tilde{Pr}(y)\penalty 10000\ \penalty 10000\ \forall y\end{array}$ (6)
However, without feature constraints we do not incorporate the known structural information into our choice. If we nevertheless solve this program and again apply the Principle of Maximum Entropy using the resulting $\tilde{P}(W)$, there is no guarantee that the found $P(W)$ which satisfies equation 1 will also satisfy equation 4, except in the special case that $\tilde{P}(W)=P(W)$. With this insight, we further develop this approach towards ensuring that all constraints are satisfied.
4 Uncertain Maximum Entropy
Let the function $M_{\phi}\colon\tilde{P}(W)\to P_{\lambda,\phi}(W)$ be the result of solving program 2 for a given $\tilde{P}(W)$. Consider the set of all solutions $P_{\centerdot,\phi}(W)$ given all possible $\tilde{P}(W)$; as program 2 is convex, $P_{\centerdot,\phi}(W)$ must be convex and contains all distributions that satisfy equation 1 each having the maximum entropy of all other $P(W)$ with equal $\Phi$. By additionally requiring that $\tilde{P}(W)$ is in this set, we ensure all constraints are satisfied while incorporating both the information we have about the structure of $P_{0}(W)$ and that received over the communications channel.
Let us take an example. Suppose for $|\mathcal{W}|=3$, $|\mathcal{Y}|=2$ we have the following functions:
$P(Y|W)=$ $w_{1}$ $w_{2}$ $w_{3}$ $y_{1}$ 0.95 0.55 0.95 $y_{2}$ 0.05 0.45 0.05 $\tilde{P}(Y)$ = (0.9, 0.1) $\phi_{1}(W)$ = (1, 0, 0)
Figure 2: $P(W)$ probability simplex for the system described above and possible solutions: program 6 (green solid dot), applying program 2 to the results of program 6 (solid blue dot), two possible solutions using equation 5 and a uniform prior (triangles) or using the true distribution as the prior (darker triangle), and the distribution that satisfies all constraints (unfilled blue circle). For reference, the distribution with the maximum possible entropy is marked with a star, solid blue vertical line is all possible maximum entropy distributions given $\phi$, and the dashed green line is all solutions to equation 4.
All possible maximum entropy solutions $P_{\centerdot,\phi}(W)$ for the single feature function are the distributions that give equal probability mass to $w_{2}$ and $w_{3}$. We illustrate this as the vertical blue line in the probability simplex in figure 2. All $\tilde{P}(W)$ that satisfy equation 4 are shown as the dashed green line and we know that $P_{0}(W)$ must be in this set. Notice the point at the intersection of the two lines marked with an unfilled circle is the only distribution that satisfies both program 2 and equation 4, and therefore must be $P_{0}(W)$.
While in this example there is only a single satisfying solution, in general it can be the case that the system is underdetermined and therefore more than one equally satisfying distribution exists. As we still desire a best estimate of $\lambda$ we must arrive at a single distribution within this feasible set. To this end we propose again applying the Principle of Maximum Entropy to choose the one satisfying solution with the maximum entropy. Intuitively we define this new Principle of Uncertain Maximum Entropy as when choosing among multiple equally satisfying distributions that are in the set of all maximum entropy $P(W)$ with respect to $\phi$ and the set of all possible $\tilde{P}(W)$, the best choice is the one with the maximum entropy. The Principle of Uncertain Minimum Cross Entropy is similarly defined.
This new principle presented as a non-linear, hierarchical convex program with feature expectation and discrete memoryless communication channel constraints is given by:
$\begin{array}[]{l}\max-\sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\log\tilde{Pr}(w)\\ \mbox{{\bf subject to}}\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)=1\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\sum\nolimits_{x\in\mathcal{X}}Pr(x|w)Pr(y|x)=\tilde{Pr}(y)\penalty 10000\ \penalty 10000\ \forall y\\ \tilde{P}(W)=M_{\phi}(\tilde{P}(W))\end{array}$ (7)
5 Analysis and Interpretation
With the goal of justifying our claims in sections 3 and 4, we begin our analysis by showing that the Principle of Uncertain Maximum Entropy as shown in program 7 is convex.
Theorem 1.
The satisfying set of the program 7 is convex.
Proof Equation 4 is linear in $\tilde{P}(W)$ and the function $M_{\phi}(\tilde{P}(W))$, being the result of a convex program, is a convex set. As the intersection of a convex set and a linear set must be convex, the satisfying set is convex.
Theorem 2.
Program 7 is convex.
Proof The proof is by contradiction. Suppose we claim that two distributions $P(W)_{1}$ and $P(W)_{2}$ satisfy all constraints while $P(W)_{1}$ and $P(W)_{2}$ both have the maximum entropy. Let $0\leq\alpha\leq 1$ and define $P(W)_{\alpha}=\alpha P(W)_{1}+(1-\alpha)P(W)_{2}$. Define the entropy of $P(W)_{\alpha}$ as a function of $\alpha$, $H(\alpha)$, where $H(1)=H(P(W)_{1})$ and $H(0)=H(P(W)_{2})$.
$H(\alpha)$ $=H(\alpha P(W)_{1}+(1-\alpha)P(W)_{2})$
$\frac{d^{2}}{d\alpha^{2}}H(\alpha)$ $=\frac{d^{2}}{d\alpha^{2}}H(\alpha P(W)_{1}+(1-\alpha)P(W)_{2})(-P(W)_{1}+P(W)_{2})^{2}$
$=\frac{d^{2}}{d\alpha^{2}}H(P(W)_{\alpha})(-P(W)_{1}+P(W)_{2})^{2}$
As entropy is a convex function, $\frac{d^{2}}{d\alpha^{2}}H(P(W)_{\alpha})<0$. So if $P(W)_{1}\neq P(W)_{2}$ then $H(\alpha)$ is a convex function in $\alpha$ as $\frac{d^{2}}{d\alpha^{2}}H(\alpha)<0$. But then at least one of $P(W)_{1}$ and $P(W)_{2}$ cannot be the maximal point.
Corollary 2.1.
The solution to the principle of uncertain maximum entropy is unique.
Since in a convex program any local optima must be the global optimum the Principle of Uncertain Maximum Entropy may be thought of as a function $U_{\phi,P(Y|W)}\colon\tilde{P}(Y)\to P_{\lambda,\phi}(W)$.
Corollary 2.2.
No other satisfying $P(W)$ exists with more entropy than the unique solution.
Given that entropy is the opposite of information, we conclude that any other satisfying choice contains more information than the one with maximum entropy. Given that a higher entropy equally satisfying alternative exists, the added information from another choice can not be justifiably attributed to the constraints, which represent all knowledge of $P_{0}(W)$. Therefore, this added information must be a form of bias. All else being equal, the solution with the least bias is expected to be least wrong most often. These facts allow us to arrive at the conclusion that justifies the principle; any other satisfying choice is expected to be worse averaged over all scenarios than that with the maximum entropy $P(W)$.
The Principle of Uncertain Maximum Entropy generalizes the classic principle and the Principle of Latent Maximum Entropy (see Appendix A).
Theorem 3.
The Principle of Maximum Entropy is a special case of the Principle of Uncertain Maximum Entropy in which only one $\tilde{P}(W)$ satisfies equation 4.
Proof In the event that only one $\tilde{P}(W)$ is possible given $\tilde{P}(Y)$ we may reduce program 7 to only $P(W)=M_{\phi}(\tilde{P}(W))$ as the removed constraints and objective are trivially satisfied. This is identical to program 2.
As $P_{0}(W)$ must be located within the feasible set and the Principle of Uncertain Maximum Entropy dictates we choose the member of this set with the maximum entropy, we must have that $H(P_{0}(W))\leq H(U_{\phi,P(Y|W)}(\tilde{P}(Y)))$. Therefore, some amount of information may be inaccessible to the receiver and we may consider this information lost to the ”environment”, either through the encoding or transmission, when such a communication channel is in use. Let $E_{\phi}(W;Y)=H(U_{\phi,P(Y|W)}(\tilde{P}(Y)))-H(P_{0}(W))$ be the amount of information lost in this way. Without knowing $H(P_{0}(W))$ we cannot determine how much information is lost, however, we may upper bound $E_{\phi}(W;Y)$ by finding the minimal entropy of any $P(W)$ within the satisfying set and then $E_{\phi}(W;Y)\leq H(U_{\phi,P(Y|W)}(\tilde{P}(Y)))-\min H$. We may now see that the Principle of Uncertain Maximum Entropy produces an upper bound on $H(P_{0}(W))$, or equivalently, a lower bound on its information content.
6 Multi-Channel Communications
The principle of uncertain maximum entropy readily generalizes to multi-channel scenarios. Suppose a given message’s encoding is received simultaneously over more than one channel by the receiver. We simply consider the joint signal distribution as: $\mathcal{Y}=\mathcal{Y}_{1}\cup\mathcal{Y}_{2}$ and alter $P(Y|X)$ accordingly, for instance if $\mathcal{Y}_{1}$ is independent of $\mathcal{Y}_{2}$ then $P(Y|X)\penalty 10000\ =\penalty 10000\ P(Y_{1}|X)P(Y_{2}|X)$.
Alternatively, if separate messages sampled from the same $P_{0}(W)$ are sent over unique channels, we simply add constraints for each unique encoding and channel in use. Let $\mathbb{C}$ be the set of all communication channel tuples $c$: ($\mathcal{X},\mathcal{Y},P(X|W),P(Y|X),\tilde{P}(Y)$). Now for each $c\in\mathbb{C}$ we modify program 4 as follows:
[…]
7 Solution
Due to the presence of the convex function $M_{\phi}(\tilde{P}(W))$ in program 7 bilevel optimization techniques are required to solve it. We present a simple algorithm based upon decomposition (Boyd and Vandenberghe, 2004) in algorithm 1.
Algorithm 1 uMaxEnt
1:Start: Initialize $Pr(w)$ to $\log|\mathcal{W}|\penalty 10000\ \forall\penalty 10000\ w$
2:Solve the following convex program to arrive at a new $\tilde{P}(W)$:
$\begin{array}[]{l}\min\sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\log\frac{\tilde{Pr}(w)}{Pr(w)}\\ \mbox{{\bf subject to}}\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)=1\\ \sum\nolimits_{w\in\mathcal{W}}\tilde{Pr}(w)\sum\nolimits_{x\in\mathcal{X}}Pr(x|w)Pr(y|x)=\tilde{Pr}(y)\penalty 10000\ \penalty 10000\ \forall y\end{array}$ (13)
3:Solve program 2 to arrive at a new $P(W)$.
4:Repeat at step 1: Until convergence of $P(W)$
5:Output: $P(W)$
Notice that after step 2, $P(W)$ will either be equal to or have less entropy than the previous $P(W)$. Then, the $\tilde{P}(W)$ found on the next iteration must have equal or less entropy than the current $\tilde{P}(W)$. Therefore, the algorithm must eventually terminate. By beginning at the maximum entropy distribution possible, we ensure that the convergence occurs at the solution to program 7. In the event a prior is given other than uniform, we set $P(W)$ to the prior at start and change step 2 to solve the Principle of Minimum Cross Entropy.
8 Experimental Performance
We experimentally compare the performance of the Principle of Uncertain Maximum Entropy to a number of existing techniques in randomly generated programs. We sample from $P_{0}(W)$ NN times (where $N$ may be infinite), then with each $w$ we sample $x\sim P(X|W=w)$ and $y\sim P(Y|X=x)$ respectively, these samples are used to generate $\tilde{P}(W)$ and $\tilde{P}(Y)$ as needed. We use the L-BFGS (Liu and Nocedal, 1989) algorithm to solve the Lagrangian dual of all programs. The algorithms we examine are:
  • uMaxEnt - Is performed as described in algorithm 1 using $\tilde{P}(Y)$.
  • MaxEnt - Program 2 is solved using $\tilde{P}(W)$, providing a best-case control.
  • mlMaxEnt - $\tilde{P}(W)$ is first computed using the most likely $w$ in equation 5 for each $y$ sample, then program 2 is solved.
  • dMaxEnt - $\tilde{P}(W)$ is computed according to equation 6 given $\tilde{P}(Y)$, then program 2 is solved.
Where indicated we consider the Kullback-Lieblier divergence $D_{KL}(P_{\lambda,\phi}(W)\mathrel{\|}P_{0}(W))$ as our metric of accuracy. We use $|\mathcal{W}|=10$ for all experiments. More details are available in Appendix B.
We first note that all algorithms are equivalent when $W=X=Y$, and in our tests all achieve an average $D_{KL}$ $\approx$ 0 as expected. In the event that $|\mathcal{Y}|=|\mathcal{W}|$ but $Y\neq W$ then mlMaxEnt dramatically reduces in performance relative to all other algorithms, as can be seen in table 1. uMaxEnt and dMaxEnt both produce low average error, however, these averages are significantly affected by the multi-modal nature of our results. We observe two error modes in our results, where over all experimental configurations except $Y=W$, 4242% achieve a $D_{KL}\geq$ $0.001$ and 5353% achieve $D_{KL}\leq$ 0.000 000 10.000\,000\,1. The primary factors determining the performance of uMaxEnt are the number of features $|\phi|$ and the size of the signal alphabet $|\mathcal{Y}|$.
MaxEnt uMaxEnt dMaxEnt mlMaxEnt
$Y=W$ 0.000 000 000 000 003 20.000\,000\,000\,000\,003\,2 0.000 000 000 000 003 10.000\,000\,000\,000\,003\,1 0.000 000 000 000 001 60.000\,000\,000\,000\,001\,6 0.000 000 000 000 001 40.000\,000\,000\,000\,001\,4
$|\mathcal{Y}|=|\mathcal{W}|$ 0.000 000 000 000 4390.000\,000\,000\,000\,439 0.001814 0.01824 $1.0398$
Table 1: Average $D_{KL}$ of the experimentally tested algorithms when $|\mathcal{Y}|=|\mathcal{W}|$, $|\phi|$ varies from 1 to 9.
Figure 3: Box plot of the $D_{KL}$ achieved by uMaxEnt, dMaxEnt, and Multi-Channel uMaxEnt for two channels as $|\phi|$ is varied. Points mark the respective mean, median is the line.
Figure 4: Box plot of the $D_{KL}$ achieved by uMaxEnt, dMaxEnt, and Multi-Channel uMaxEnt for two channels as $|\mathcal{Y}|$ is varied. Points mark the respective mean, median is the line.
Figure 5: Scatter plot of the $D_{KL}$ achieved by multi-channel uMaxEnt against single-channel uMaxEnt, 20,000 points randomly selected from our data set (where $Y\neq W$). Note log scales on X and Y axes.
As can be seen in figure 3, the smaller the number of features the lower the average $D_{KL}$ achieved, with uMaxEnt always lower than dMaxEnt. Importantly, the inter-quartile range in figure 3 shows uMaxEnt dramatically outperforms dMaxEnt when features are below 5, with median $D_{KL}$ values many orders of magnitude smaller. This is attributed to the more tightly constrained feasible set producing a smaller area for $P_{0}(W)$ to be in, which uMaxEnt exploits, finding solutions with less entropy than dMaxEnt. Further examining the inter-quartile ranges, below 5 features most solutions are in the low error mode of our results, and above 5 features most solutions are in the higher error mode. By contrast, dMaxEnt is not as significantly affected by the number of features, with both the average and median $D_{KL}$ reducing by about one order of magnitude as the features go from 9 to 1.
Because uMaxEnt is producing solutions with a minimal amount of bias, we may attribute the reduction in entropy realtive to dMaxEnt to the information contained in the features. However, as we show in figure 6 where we compare uMaxEnt against the feature-free program 6, we cannot consider the information gain from features separate from $Y$. Notice that the information increase is usually highest when $|\phi|=|\mathcal{Y}|-1$, when they have the same degrees of freedom.
Figure 4 examines the performance of our algorithms as $|\mathcal{Y}|$ is varied. Below $|\mathcal{Y}|=6$, most solutions are in the high error mode of our results and above 6 in the low error mode. Both uMaxEnt and dMaxEnt achieve a similar median $D_{KL}$ when $|\mathcal{Y}|$ is 10, however, examining the inter-quartile range shows uMaxEnt is significantly more consistent. In this case, the solutions that cause uMaxEnt to have an average $D_{KL}$ so significantly higher than the 3rd quartile are almost entirely those with $|\phi|\geq 7$, which is not also the case for dMaxEnt. Overall, the smaller $|\mathcal{Y}|$ is, the less information is available in a given communications channel and we see an expected increase in error in our results.
Figure 6: Information increase in the solutions found by uMaxEnt relative to that found by program 6 alone for varying sizes of $|\phi|$ and $|\mathcal{Y}|$, $Y\neq W$.
8.1 Multi-Channel uMaxEnt
Due to the loss of information when $\mathcal{Y}\neq\mathcal{W}$, we cannot expect uMaxEnt to produce low error solutions for all possible communication channels. However, by adding additional channels the feasible set can be constrained further leading to lower error. We show this in figures 3 and 4 as the multi-channel uMaxEnt with just one additional channel significantly outperforms single-channel uMaxEnt for all feature counts and $|\mathcal{Y}|$. We visualize this increase in performance further in figure 5, notice particularly the top-left cluster; these points have a high $D_{KL}$ found by single-channel uMaxEnt but with just one extra channel the $D_{KL}$ has dropped dramatically. The top-right cluster shows that some high $D_{KL}$ results are still present and whose error may be improved upon with additional channels.
8.2 uMaxEnt Approximation
uMaxEnt requires that $\tilde{P}(Y)$ is estimated to arbitrary accuracy and therefore that $\tilde{P}(W)=P_{0}(W)$. However, this is only guaranteed in scenarios where $\tilde{P}(Y)$ is perfectly known or the number of samples is infinite. For all other scenarios we must relax our requirements appropriately, and we present two simple modifications to utilize uMaxEnt with finite $N$.
  • As $\tilde{P}(W)$ and $P_{0}(W)$ are not necessarily equal, we relax step 3 such that we repeat while $D_{KL}(\tilde{P}(W)\mathrel{\|}P(W))>\epsilon$, where $\epsilon$ is some maximum tolerance on the divergence between these two distributions.
  • As $\tilde{P}(Y)$ may be poorly estimated, satisfying $\tilde{P}(W)$ may be out of the probability simplex. We therefore add a Gaussian regularization term $\sum\nolimits_{k}\lambda_{k}^{2}/N$ to the Lagrangian dual of step 1.
8.2.1 Upper Bound on $\epsilon$ from the Law of Large Numbers
We may incorporate additional information we have about the relationship between $\tilde{P}(W)$ and $P_{0}(W)$ by choosing $\epsilon$ appropriately. While ad-hoc choices based upon information about the sampling process are possible, here we present a broadly applicable bound on $\epsilon$ based upon the strong law of large numbers (Cover and Thomas, 2006) that depends only on the assumption that $\tilde{P}(W)$ is produced by i.i.d. sampling N times from $P_{0}(W)$.
The strong law of large numbers is:
$Pr(D_{KL}(\tilde{P}(W)\mathrel{\|}P_{0}(W))>\epsilon)\leq 2^{-N(\epsilon-|\mathcal{W}|\frac{\log_{2}(N+1)}{N})}$
Which states that as $N\rightarrow\infty$ the divergence of the empirical distribution from the true distribution goes to zero with probability 1. We make use of this result to find an upper bound for $\epsilon$. Bounding the above inequality by 1 and solving for $\epsilon$, we find:
$\epsilon\leq\frac{|\mathcal{W}|log_{2}(N+1)}{N}$ (14)
Notice $\epsilon\rightarrow 0$ as $N\rightarrow\infty$. Thus, for large $N$, this bound on $\epsilon$ ensures that any $P(W)$ output by uMaxEnt must have low divergence to the empirical distribution.
8.2.2 Performance of the Approximation
Figure 7: Average $D_{KL}$ achieved by varying algorithms as the number of samples increases. Performance at infinity is marked with a point. Note log scale on the X axis.
We show the typical performance of our approximation in figure 7 for $|\mathcal{Y}|=5,\alpha=3,|\phi|=(1,2,3,4)$. Notice that uMaxEnt begins to show a reduction in average $D_{KL}$ around $N$ = 500. The performance of non-regularized MaxEnt and mlMaxEnt are shown for reference. Notice the points showing performance at infinity, indicating that uMaxEnt will still significantly improve with further samples while dMaxEnt is near its maximum performance at $N=10^{6}$.
Note that due to the approximate nature of this algorithm, it is difficult to provide firm performance guarantees as outliers may be significant. For instance for the specific case $|\mathcal{Y}|=3,|\phi|=2,$ and $N=65536$ the average $D_{KL}$ of uMaxEnt = 0.3721 while dMaxEnt achieves 0.3267. Despite this, we observe in our experiments that the median $D_{KL}$ of uMaxEnt is always less than or equal that of dMaxEnt.
9 Discussion
In our experiments the use of one additional communication channel showed a reduction in the average $D_{KL}$ and therefore less information loss on average. Ignoring the special case where $W=Y$, as we continue adding random channels to $\mathbb{C}$ we expect that the set of $P_{0}(W)$ for which more than one $\tilde{P}(W)$ satisfies the constraints goes to zero. Therefore, we may interpret the classic Principle of Maximum Entropy as an infinite-channel Principle of Uncertain Maximum Entropy, where we arrive at a solution with no possible information lost to the environment and the entropy is minimal with no way of increasing information further without adding bias.
In this work we assume that both $\phi$ and $P(Y|W)$ are known to the receiver. Future directions include relaxing these assumptions, for instance, if $\phi$ is not completely known $P_{0}(W)$ is expected to have un-modeled structural variables or otherwise some amount of noise. In this situation, uMaxEnt must be modified as $P_{0}(W)$ may not lie within the $P_{\centerdot,\phi}(W)$.
As an example, suppose we have some bound on the maximum divergence between $P_{\centerdot,\phi}(W)$ and $P_{0}(W)$, call it $\epsilon_{M}$. We could alter uMaxEnt such that after step 2 we find the point $a$ along the line segment between $\tilde{P}(W)$ and $P(W)$ such that $D_{KL}(a\mathrel{\|}P(W))=\epsilon_{M}$. Then set $P(W)=a$ and continue, outputting $M_{\phi}(P(W))$ at the end. This will ensure we find the point consistent with observations and the maximum allowed divergence from $P_{\centerdot,\phi}(W)$.
Approximate performance of uMaxEnt was shown for large but limited $N$ but no increase for smaller $N$ over dMaxEnt was observed. This indicates that $\epsilon$ is large enough that only one iteration of uMaxent is performed, therefore performance improvement may be possible with tighter bounds on $\epsilon$.
Acknowledgments and Disclosure of Funding
This work was partially funded by NSF grant 1830421. The authors wish to thank our colleagues in the computer science department at UNCA for their support and comments.
Appendix A
The Principle of Latent Maximum Entropy as presented by Wang et al. (2012) is a generalization of the Principle of Maximum Entropy in which some portion of each message variable $w$ is missing to the receiver and the rest is received without error. Divide each $w$ into $y$ and $z$ such that $w=y\cup z$ and $P(W)=P(Y,Z)$. $y$ is the portion of $w$ that is perfectly received while $z$ is perfectly hidden, and let $\mathcal{Z}_{y}$ be the set of all $z$ which when combined with a given $y$ forms a $w\in\mathcal{W}$. The principle of latent maximum entropy is now defined as:
$\max\left(-\sum\nolimits_{w\in\mathcal{W}}Pr(w)\log Pr(w)\right)$
subject to
$\sum\nolimits_{w\in\mathcal{W}}Pr(w)=1$
$\sum\nolimits_{w\in\mathcal{W}}Pr(w)\phi_{k}(w)=\sum\nolimits_{y\in\mathcal{Y}}\tilde{Pr}(y)\sum\nolimits_{z\in\mathcal{Z}_{y}}Pr(z|y)\phi_{k}(w)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall k$ (15)
The authors give an Expectation-Maximization (McLachlan and Krishnan, 2007) based algorithm, EM-IS, to find solutions to program 15. The relevant steps of this algorithm are:
  • E Step: Compute $\tilde{\Phi}=\sum\nolimits_{y\in\mathcal{Y}}\tilde{Pr}(y)\sum\nolimits_{z\in\mathcal{Z}_{y}}Pr(z|y)\phi_{k}(w)\penalty 10000\ \penalty 10000\ \forall k$
  • M Step: Solve program 2
Theorem 4.
The Principle of Latent Maximum Entropy as solved with the expectation-maximization based EM-IS algorithm with an initial $Pr(w)=log|\mathcal{W}|$ is a special case of the Principle of Uncertain Maximum Entropy in which $|\mathcal{Y}|<|\mathcal{W}|$ and $P(Y|W)$ is deterministic.
Proof As $Pr(y,z)=Pr(w)$ we may rewrite the constraint of equation 15 as:
$\sum\nolimits_{w\in\mathcal{W}}Pr(w)\phi_{k}(w)=\sum\nolimits_{y\in\mathcal{Y}}\frac{\tilde{Pr}(y)}{Pr(y)}\sum\nolimits_{z\in\mathcal{Z}_{y}}Pr(w)\phi_{k}(w)$
Where $Pr(y)=\sum\nolimits_{z\in\mathcal{Z}_{y}}Pr(w)$. Let $Pr(y|w)=1$ if $\exists z\in\mathcal{Z}_{y}\colon w=(y\cup z)$, otherwise 0. This constraint is satisfied whenever $\tilde{Pr}(y)=Pr(y)=\sum\nolimits_{w\in\mathcal{W}}Pr(w)P(y|w)\penalty 10000\ \penalty 10000\ \forall y$. Therefore, program 15 as given is equivalent to program 6.
Now notice that the EM-IS algorithm restricts the search space of $P(W)$ to $P_{\centerdot,\phi}(W)$. As EM monotonically increases towards convergence, when the EM algorithm converges the found $P_{\lambda,\phi}(W)$ will satisfy the constraints of program 7. Because the search space is convex it will not converge early. Beginning the search at the maximum entropy $P(W)$ possible completes the proof.
Appendix B
Additional Experimental Details
To directly compare the performance of the four algorithms considered, we give each one identical input, or as close to identical as possible. For instance, if $P_{0}(Y)$ is given to uMaxEnt, we give the corresponding $P_{0}(W)$ used to generate it to MaxEnt. We describe the steps to accomplish this below.
B.1 Model Generation
We first fix $|\mathcal{W}|=10$ and select $|\phi|\in\{1,2,\dots,9\}$. We generate sparse feature sets by choosing a probability $p\in\{0.25,0.33,0.66\}$ and then assigning $\phi_{k}(w)=1$ with probability $p$ for each $k,w$, otherwise 0. If a given $\phi_{k}=0\penalty 10000\ \forall\penalty 10000\ w$, we re-generate this feature until at least one $w$ is assigned a $1$. We choose true weights $\lambda_{k}=U(-1,1)*\alpha$, where $\alpha$ is 2 or 3, now each $Pr_{0}(w)\propto e^{\sum\limits_{k}\lambda_{k}\phi_{k}(w)}$. We then sample $P_{0}(W)$ 1,048,576 times to provide samples needed for the approximate experiments.
B.2 Channel Generation
Given the above model $W$ and samples, choose values $\alpha_{x}\in\{DET,0.5,1,2\}$ and $\alpha_{y}\in\{DET,0.5,1,2\}$. If $\alpha_{x}=DET$, set $X=W$ and if $\alpha_{y}=DET$, set $Y=X$. Otherwise, choose $|X|\in\{5,8,10\}$ and then $|Y|\in\{2,3,\dots,|X|\}$. Now let $Pr(x|w)\propto e^{\alpha_{x}U(0,1)}\penalty 10000\ \forall\penalty 10000\ w$, $Pr(y|x)\propto e^{\alpha_{y}U(0,1)}\penalty 10000\ \forall\penalty 10000\ x$ respectively.
Using these distributions and $P_{0}(W)$, we calculate $P_{0}(X)$ and $P_{0}(Y)$ which is used in the exact experiments. For the approximate experiments, we generate 1,048,576 samples of $X$ and $Y$ by sampling $x_{i}\sim P(X|w_{i})$ and $y_{i}\sim P(Y|x_{i})$ respectively.
This process is repeated a second time for the multi-channel experiments, except we hold $|X|$ and $|Y|$ the same, and sample $\alpha_{x}$ and $\alpha_{y}$ from $U(0,2)$.
B.3 Procedure
A total of 77,760 experimental configurations were generated, distributed approximated equal among all the parameter permutations. Experiments proceeded first by loading a given model and communication channel(s), then generating $\tilde{P}(W)$ and $\tilde{P}(Y)$ using the first $N$ samples of each respectively, or setting $\tilde{P}(W)=P_{0}(W)$ and $\tilde{P}(Y)=P_{0}(Y)$ for exact experiments. Each algorithm is then input its appropriate distribution ($\tilde{P}(W)$ for MaxEnt, $\tilde{P}(Y)$ all others) and solved. The resulting $P_{\lambda,\phi}(W)$’s were recorded and $D_{KL}(P_{\lambda,\phi}(W)\mathrel{\|}P_{0}(W))$ was calculated.
We use the NLOpt (Johnson, 2007) library version v2.10.0 to solve our programs, specifically the Lagrangian Dual of each program is solved with L-BFGS (Liu and Nocedal, 1989). Tolerances were set as low as possible to ensure accuracy while avoiding numerical instability, with change in parameter tolerance 0.000 000 000 000 010.000\,000\,000\,000\,01, maximum evaluations 100 000100\,000, and absolute maximum parameter value of $1000$.
For uMaxEnt, iterations were stopped after a maximum of $2000$, or when the $D_{KL}(\tilde{P}(W)\mathrel{\|}P(W))\leq\epsilon$. For approximate experiments, $\epsilon$ is set to the value computed from section 8.2, for exact experiments $\epsilon=$ 0.000 000 000 010.000\,000\,000\,01.